# Understanding and Mitigating Compositional Issues in Text-to-Image Generative Models

### v1

@@ 4 Source (ii) : Sub-optimality of CLIP Text-Encoder for Compositional Prompts
In this section, we understand if the UNet is capable of generating compositional scenes by optimizing the text-embeddings that it takes as the conditional input. Given an input prompt $c$ with a particular composition (e.g., “a red book and a yellow table”), we utilize our dataset and obtain $\mathcal{D}_{c}$ including high-quality compositional images for prompt $c$. We then optimize the output text-embedding $\mathbf{c}$ as follows:
@@
$\mathbf{c}^{*}=\arg\min_{\mathbf{c}}\mathbb{E}_{x_{0}\sim\mathcal{D}_{c},\epsilon,t}\left[\left\lVert\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t)\right\rVert_{2}^{2}\right].$
@@
We then use $\mathbf{c}^{*}$ to generate images using the UNet $\epsilon_{\theta}$ across different seeds. Figure 3 depicts a few of generated images using optimized text-embeddings.
@@
Key Results. As seen in Figure 3, we consistently improve VQA scores across a variety of compositional prompts (i.e., color, texture, and shape). This indicates that CLIP text-encoder does not output the proper text-embedding suitable for generating compositional scenes. However, that optimized embedding space exists, highlighting the ability of UNet to generate coherent compositional scenes when a proper text-embedding is given. This further motivates the idea of improving CLIP output space to mitigate compositionality issues in text-to-image diffusion models. We refer to Appendix B for other configurations of optimizing text-embeddings where we observe that optimizing only a subset of few tokens, can also effectively improve compositionality.
5 Linear Projection on CLIP: A Simple Baseline to Improve Compositionality in Text-to-Image Generative Models
In this Section, we provide two baselines CLP and WiCLP that are linear modification of CLIP output to map that sub-optimal space to an enhanced one, better suited for compositionality.
@@ 5.1 CLP: Token-wise Compositional Linear Projection
Given the text-embedding $\mathbf{c}\in\mathbb{R}^{n\times d}$ as the output of the text-encoder for prompt $c$, i.e., $\mathbf{c}=v_{\gamma}(c)$, we train a linear projection $\texttt{{CLP}}_{W,b}:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times d}$. This projection includes a matrix $W\in\mathbb{R}^{d\times d}$ and a bias term $b\in\mathbb{R}^{d}$, which are applied token-wise to the output text-embeddings of the encoder. More formally, for $\mathbf{c}\in\mathbb{R}^{n\times d}$ including text-embeddings of $n$ tokens $\mathbf{c}_{1},\mathbf{c}_{2},\dotsm,\mathbf{c}_{n}\in\mathbb{R}^{d}$, $\texttt{{CLP}}_{W,b}(\mathbf{c})$ is obtained by stacking projected embeddings $\mathbf{c}^{\prime}_{1},\mathbf{c}^{\prime}_{2},\dotsm,\mathbf{c}^{\prime}_{n}$ where $\mathbf{c}^{\prime}_{i}=W^{T}\mathbf{c}_{i}+b$.
@@
$W^{*},b^{*}=\arg\min_{W,b}\mathbb{E}_{(x_{0},c)\sim\mathcal{D},\epsilon,t}\left[\left\lVert\epsilon-\epsilon_{\theta}]\left(\mathbf{z}_{t},\texttt{{CLP}}_{W,b}\left(\mathbf{c}\right),t\right)\right\rVert_{2}^{2}\right].$
@@ 5.2 WiCLP: Window-based Compositional Linear Projection
In this Section, we propose a more advanced linear projection scheme where the new embedding of a token is derived by applying a linear projection on that token in conjunction with a set of its adjacent tokens, i.e., tokens within a specified window. This method not only leverages the benefits of CLP but also incorporates the contextual information from neighboring tokens, potentially leading to more precise text-embeddings.
@@
$W^{*},b^{*}=\arg\min_{W,b}\mathbb{E}_{(x_{0},c)\sim\mathcal{D},\epsilon,t}\left[\left\lVert\epsilon-\epsilon_{\theta}]\left(\mathbf{z}_{t},\texttt{{WiCLP}}_{W,b}\left(\mathbf{c}\right),t\right)\right\rVert_{2}^{2}\right].$
Note that we use $s=2$, i.e., window length of $5$ in our experiments.
prompt: ”A bathroom with green tile and a red shower curtain”
$\tau=1000$ (No Guidance) $\tau=900$ $\tau=800$ $\tau=600$ $\tau=400$ $\tau=200$
@@ 5.3 Switch-Off: Trade-off between Compositionality and Clean Accuracy
Fine-tuning models or adding supplementary modules to a base model often results in a degradation of image quality and an increase in the Fréchet Inception Distance (FID) score. To balance the trade-off between improved compositionality and the quality of generated images for clean prompts – an important issue in existing work – inspired by Hertz et al. [10], we adopt Switch-Off, where we apply the linear projection only during the initial steps of inference. Specifically, given a time-step threshold $\tau$, for $t\geq\tau$, we use $\texttt{{WiCLP}}_{W^{*},b^{*}}(\mathbf{c})$, while for $t<\tau$, we use the unchanged embedding $\mathbf{c}$ as the input to the cross-attention layers.
@@
Figure 5 illustrates the trade-off between VQA score and FID on a randomly sampled subset of MS-COCO [14] for different choices of $\tau$. As shown, even a large value of $\tau$ suffices for obtaining high-quality compositional scenes as the composition of final generated image is primarily formed at early steps. Thus, choosing a large $\tau$ preserves the model’s improved compositionality while maintaining its clean accuracy. Setting $\tau=800$ offers a competitive VQA score compared to the model where projection is applied at all time steps, and achieves a competitive FID similar to that of the clean model. Figure 4 depicts a few images generated using different choices of $\tau$. We refer to Appendix G for more visualizations.
5.4 Experiments
@@
Training Setup. All of the models are trained using the objective function of diffusion models on color, texture, and shape datasets. During training, we keep all major components frozen, including the U-Net, CLIP text-encoder, and VAE encoder and decoder, and only the linear projections are trained. We refer to Appendix D for details on training procedure.
Improved cross-attention maps. Figure 6 illustrates cross-attention maps for a sample prompt. In the baseline model, the attention maps are flawed, with some tokens incorrectly attending to the wrong pixels. However, with both CLP and WiCLP, objects and attributes more accurately attend to their respective pixels. For more visualizations, see Appendix F.
Results. Figure 2 presents images generated when applying CLP and WiCLP. When generating compositional prompts with a baseline model, objects are often missing or attributes are incorrectly applied. However, with CLP and WiCLP, objects and their corresponding attributes are more accurately generated. We refer to Appendix E for more visualizations. We compare VQA scores of our proposed methods and other state-of-the-art methods in Table 1. As shown, both CLP and WiCLP significantly improve upon the baselines. WiCLP achieves higher VQA scores compared to other state-of-the-art methods, despite its simplicity. For evaluation, we relied on the VQA score, identified by [12] as the most informative metric. While GORS [12] achieves a comparable VQA score, it fine-tunes the entire model, whereas our method uses a simple linear projection with about 200 times fewer parameters. Additionally, our approach yields lower FID scores, highlighting another advantage. More precisely, the FID scores for the models are: GORS ($30.54$), SD v1.4 ($24.33$), SD v2 ($23.27$), WiCLP on top of SD v1.4 ($25.40$), and WiCLP on top SD v2 ($27.40$) when using Switch-Off with $\tau=800$.
Comparison between CLP and WiCLP. We observe that WiCLP improves over CLP (special case of WiCLP with $s=0$) by incorporating adjacent tokens in addition to the actual token. This approach likely improves embeddings by mitigating unintended attention from adjacent tokens. For discussion on choosing the window length ($s$) in WiCLP, see Appendix H.
Model Color Texture Shape
Stable Diffusion v1.4 [21] Baseline 0.3765 0.4156 0.3576
[…]
WiCLP 0.5383 0.5671 0.4527
Stable Diffusion v2 [21] Baseline 0.5065 0.4922 0.4221
Composable [16] 0.4063 0.3645 0.3299
Structured [8] 0.4990 0.4900 0.4218
Attn-Exct [4] 0.6400 0.5963 0.4517
GORS-unbaised [12] 0.6414 0.6025 0.4546
[…]
WiCLP 0.6500 0.6036 0.4821
@@
Figure 6: Applying CLP results in more accurate cross-attention maps.
@@ 6 Conclusion
Our paper examines potential error sources in text-to-image models for generating images from compositional prompts. We identify two error sources: (i) A minor error source, where the token embeddings in the CLIP text-encoder have erroneous attention contributions and (ii) A major error source, where we find the output space of the CLIP text-encoder to be sub-optimally aligned to the UNet for compositional prompts. Leveraging our observations, we propose a simple and strong baseline WiCLP which involves fine-tuning a linear projection on CLIP’s representation space. WiCLP though inherently simple and parameter efficient, outperforms existing methods on compositional image generation benchmarks and maintains a low FID score on a broader range of clean prompts. We discuss limitations in Appendix A.

### v2

@@ 3.2 Source (ii) : Sub-optimality of CLIP Text-Encoder for Compositional Prompts
In this section, we investigate whether the UNet is capable of generating better compositional scenes if provided with alternative (improved) text embeddings, rather than relying on the output of the CLIP text-encoder. For a given input prompt $p$ with a specific composition (e.g., “a red book and a yellow table”), we utilize our dataset (described in Section 8) to obtain $\mathcal{D}_{p}$, a set of high-quality compositional images corresponding to prompt $p$. Next, we extract the text embedding $\mathbf{c}$ from the CLIP text-encoder for prompt $p$. Using this embedding as the initialization, we create a learnable vector $\mathbf{c}^{*}$ of the same dimensionality. Keeping all other components (such as the UNet) frozen, we optimize this learnable vector as follows:
@@
$\mathbf{c}^{*}=\arg\min_{\mathbf{c}}\mathbb{E}_{x_{0}\sim\mathcal{D}_{p},\epsilon,t}\left[\left\lVert\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t)\right\rVert_{2}^{2}\right].$
@@
We then use the optimized text embedding $\mathbf{c}^{*}$ to generate images with the UNet $\epsilon_{\theta}$. Figure 5 illustrates the complete pipeline.
Key Results. Utilizing Stable Diffusion v1.4, we optimize optimize $\mathbf{c}^{*}$ for all compositional prompts across the color, texture, and shape categories in the T2I-CompBench dataset. By generating samples with $\mathbf{c}^{*}$ and comparing them to those generated using $\mathbf{c}$, we observe a significant improvement in the VQA scores. As shown in Figure 5, CLIP text embeddings yield VQA scores of $0.3615$ for color, $0.4306$ for texture, and $0.3619$ for shape. In contrast, the optimized embeddings achieve $0.7513$ for color, $0.7254$ for texture, and $0.5873$ for shape.
@@
These results indicate that CLIP text-encoder does not output the proper text-embedding suitable for generating compositional scenes. However, the existence of an optimized embedding space demonstrates that the UNet can generate coherent compositional outputs when provided with appropriately improved embeddings. This finding motivates the idea of improving the CLIP output space to mitigate compositionality issues in text-to-image diffusion models. For additional configurations, including results from optimizing a subset of tokens to improve compositionality, refer to Appendix A.
4 Projection Layer for Enhancing Compositionality in the CLIP Text Embedding Space
Building on our previous findings, we focus on improving the text embedding space utilized in text-to-image generative models. Specifically, we propose learning a projection layer over the CLIP output embedding space to transform its sub-optimal representation into an enhanced space better suited for compositionality. In the following sections, we introduce two methods, CLP and WiCLP, which implement linear projections of the CLIP output embedding space to achieve this enhancement.
@@ 4.1 CLP: Token-wise Compositional Linear Projection
Given the text-embedding $\mathbf{c}\in\mathbb{R}^{n\times d}$ as the output of the text-encoder for prompt $c$, i.e., $\mathbf{c}=v_{\gamma}(c)$, we train a linear projection $\texttt{{CLP}}_{W,b}:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times d}$. This projection includes a matrix $W\in\mathbb{R}^{d\times d}$ and a bias term $b\in\mathbb{R}^{d}$, which are applied token-wise to the output text-embeddings of the text-encoder. More formally, for $\mathbf{c}\in\mathbb{R}^{n\times d}$ including text-embeddings of $n$ tokens $\mathbf{c}_{1},\mathbf{c}_{2},\dotsm,\mathbf{c}_{n}\in\mathbb{R}^{d}$, $\texttt{{CLP}}_{W,b}(\mathbf{c})$ is obtained by stacking projected embeddings $\mathbf{c}^{\prime}_{1},\mathbf{c}^{\prime}_{2},\dotsm,\mathbf{c}^{\prime}_{n}$ where $\mathbf{c}^{\prime}_{i}=W^{T}\mathbf{c}_{i}+b$.
@@
$W^{*},b^{*}=\arg\min_{W,b}\mathbb{E}_{(x_{0},c)\sim\mathcal{D},\epsilon,t}\left[\Phi_{\texttt{{CLP}}}\right]$
$\Phi_{\texttt{Proj}}=\left\lVert\epsilon-\epsilon_{\theta}\left(\mathbf{z}_{t},{\scriptstyle\texttt{Proj}}_{W,b}\left(\mathbf{c}\right),t\right)\right\rVert_{2}^{2}$
@@ 4.2 WiCLP: Window-based Compositional Linear Projection
In this section, we propose a more advanced linear projection scheme where the new embedding of a token is derived by applying a linear projection on that token in conjunction with a set of its adjacent tokens within a specified window. This method not only leverages the benefits of CLP but also incorporates the contextual information from neighboring tokens, potentially leading to more precise text-embeddings.
@@
$W^{*},b^{*}=\arg\min_{W,b}\mathbb{E}_{(x_{0},c)\sim\mathcal{D},\epsilon,t}\left[\Phi_{\texttt{{WiCLP}}}\right]$
”A bathroom with green tile and a red shower curtain”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\tau=1000$ (No Guidance) $\tau=900$ $\tau=800$ $\tau=600$ $\tau=200$
@@
We observe that WiCLP improves over CLP (special case of WiCLP with $s=0$) by incorporating adjacent tokens along with the token itself. This approach enhances embeddings by reinforcing the contributions of relevant adjacent tokens. See Appendix D.8 for details on window length in WiCLP.
@@ 4.3 Switch-Off: Trade-off between Compositionality and Clean Accuracy
Fine-tuning models or adding modules to a base model often results in a degradation of image quality and an increase in the Fréchet Inception Distance (FID) score. To balance the trade-off between improved compositionality and the quality of generated images for clean prompts, nspired by Hertz et al. (2022), we adopt Switch-Off, where we apply the linear projection only during the initial steps of inference. Specifically, given a time-step threshold $\tau$, for $t\geq\tau$, we use $\texttt{{WiCLP}}_{W^{*},b^{*}}(\mathbf{c})$, while for $t<\tau$, we use the unchanged embedding $\mathbf{c}$ as the input to the cross-attention layers.
@@
Figure 7 illustrates the trade-off between VQA score and FID on a randomly sampled subset of MS-COCO (Lin et al., 2014) for different choices of $\tau$. As shown, even a large value of $\tau$ suffices for obtaining high-quality compositional scenes as the composition of final generated image is primarily formed at early steps. Thus, choosing a large $\tau$ preserves the model’s improved compositionality while maintaining its clean accuracy. Setting $\tau=800$ offers a competitive VQA score compared to the model where projection is applied at all time steps, and achieves a competitive FID similar to that of the clean model. Figure 7 depicts a few images generated using different choices of $\tau$. We refer to Appendix D.7 for more visualizations.
5 Experiments
Existing Baselines. We evaluate the performance of multiple methods alongside standard models SD v1.4, SD v2, SDXL (Podell et al., 2023), SD v3 (Esser et al., 2024), and PixArt-$\alpha$ (Chen et al., 2023). These include Composable Diffusion (Liu et al., 2022b), which addresses concept conjunction and negation in pretrained diffusion models; Structured Diffusion (Feng et al., 2022), which focuses on attribute binding; Attn-Exct (Chefer et al., 2023), which ensures correct attention to all subjects in the prompt using iterative optimizations; GORS (Huang et al., 2023), which fine-tunes Stable Diffusion v2 using a reward function; GLIGEN (Li et al., 2023), which utilizes grounding inputs such as bounding boxes; RealCompo (Zhang et al., 2024), which integrates spatial-aware diffusion models; and FLUX (Labs, 2024).
@@ 5 Experiments
Training Setup. All of the models are trained using the objective function of diffusion models on color, texture, and shape datasets described in Section 8. During training, we keep all major components frozen, including the U-Net, CLIP text-encoder, and VAE encoder and decoder, and only the linear projections are trained. We refer to Appendix D.1 for details on the training procedure.
@@
WiCLP Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: More accurate cross-attention maps using CLP.
Dataset Collection. We utilize the T2I-CompBench dataset (Huang et al., 2023), a well-recognized dataset for compositionality, focusing on three categories: color, texture, and shape, with a total of 1,000 prompts across both training and evaluation sets. To generate high-quality images, we use three generative models: SD v1.4, DeepFloyd, and SynGen (Rassin et al., 2024), creating 210 samples per prompt. This ensures a wide variety of generated images, leveraging each model’s strengths. From these, we selected the top 30 with the highest VQA scores to ensure the final dataset consists of images that best reflect the prompts.
Furthermore, for SDXL, SD v3, and PixArt-$\alpha$, we explored training WiCLP (WiCLP* in Table 1) on a higher-quality dataset generated by newer text-to-image models, such as SDXL and SD v3. Importantly, leveraging an appropriately curated dataset results in a substantial improvement in VQA scores, highlighting the importance of high-quality training data for compositional understanding.
5.1 Qualitative and Quantitative Evaluation
Qualitative Evaluation. Figure 2 presents images generated when applying WiCLP. When generating compositional prompts with a baseline model, objects may be missing or attributes are incorrectly applied. However, with WiCLP, objects and their corresponding attributes are more accurately generated. See Appendix D.5 for more visualizations.
Figure 8 illustrates cross-attention maps for a sample prompt. In the base model, attention maps are flawed, with some tokens incorrectly attending to the wrong pixels. However, with both CLP and WiCLP, objects and attributes more accurately attend to their respective pixels. For more visualizations, see Appendix D.6.
Model Color Texture Shape
SD v1.4 Baseline 0.3765 0.4156 0.3576
[…]
CLP 0.4837 0.5312 0.4307
WiCLP 0.5383 0.5671 0.4527
SD v2 Baseline 0.5065 0.4922 0.4221
Composable 0.4063 0.3645 0.3299
Structured 0.4990 0.4900 0.4218
Attn-Exct 0.6400 0.5963 0.4517
GORS 0.6414 0.6025 0.4546
[…]
CLP 0.6075 0.5707 0.4567
WiCLP 0.6500 0.6036 0.4821
SDXL Baseline 0.5770 0.5217 0.4666
[…]
WiCLP 0.6930 0.6007 0.4758
WiCLP* 0.7801 0.6557 0.5166
[…]
WiCLP*$512^{2}$ 0.5293 0.5539 0.4357
SD v3 Baseline 0.8164 0.7303 0.5852
[…]
WiCLP* 0.8213 0.7488 0.5963
Others FLUX 0.7354 0.6016 0.4777
GLIGEN 0.4288 0.3904 0.3998
RealCompo 0.7741 0.7427 0.6032
Quantitative Evaluation. Table 1 presents the VQA scores for our methods, CLP and WiCLP, alongside the baselines discussed. VQA scores of our method and other discussed baselines are provided in Table 1. As shown, both CLP and WiCLP significantly improve upon the baselines. Both methods demonstrate substantial improvements over the baselines, with WiCLP achieving the highest VQA scores among state-of-the-art approaches that utilize the same baseline model, while being more computationally and parameter-efficient. Additionally, to further validate the performance gains, we evaluated our method using additional metrics, including TIFA (Hu et al., 2023). Details of these evaluations are provided in Appendix D.2. The results demonstrated consistent improvements over the baselines, reinforcing the effectiveness of our approach. For analysis of WiCLP’s robustness and generalizability—both when trained across all categories and when applied to models using T5 text encoders—see Appendix D.3 and D.4.
Notably, our methods maintain the model’s general utility, introducing only a slight increase in the FID score; for example, experiments on MS-COCO prompts show that while our methods slightly increase FID compared to base models, this increase is smaller than that of other baselines—for instance, WiCLP achieves an FID score of $27.40$, outperforming GORS at $30.54$. Additional details on FID performance can be found in Appendix D.2.
Human Experiments. We conducted a human evaluation where participants compared images generated by SD v1.4 and SD v1.4 + WiCLP, selecting the image that best matched the given prompt. The results showed that in $34.625\%$ of cases, evaluators chose the base model; in $51.875\%$, they preferred the WiCLP; and in $13.50\%$, they rated both equally. See Appendix D.2 for further details.
5.2 Impact of WiCLP on Subsets of Tokens
To better understand the impact of WiCLP on token embeddings, we applied the trained WiCLP to specific subsets of tokens from a sample of dataset sentences. The results, shown in Fig. 10, compare the following token groups: nouns only; nouns and adjectives; nouns, adjectives, and the EOS (End of Sentence token) token; all sentence tokens; and all tokens outputted by CLIP (sentence tokens plus padding tokens). As can be seen, applying WiCLP only to a small number of tokens is sufficient for improving compositionality. Interestingly, applying WiCLP to the group of nouns, adjectives, and EOS achieves even higher VQA scores than applying WiCLP to all tokens. Despite these findings, we applied WiCLP to all tokens in our work, leaving this targeted approach for future research.
Figure 9: Effect of applying WiCLP to specific tokens. Applying WiCLP to a subset of tokens is sufficient to enhance compositionality, achieving comparable or superior performance to applying it across all tokens.
Figure 10: Performance comparison of different projection architectures on T2I-CompBench.
5.3 Alternatives to WiCLP
We explored various fine-tuning strategies for improving CLIP, including fine-tuning the entire CLIP, fine-tuning only the last layers of CLIP combined with WiCLP, and using WiCLP alone. Our results show that the original baseline model (SD v1.4) achieves a VQA score of 0.3765 on the color category of the dataset. Fine-tuning the entire CLIP without WiCLP improves the score to 0.5173, fine-tuning the last layers of CLIP combined with WiCLP achieves 0.5497, and WiCLP alone achieves 0.5383.
These findings highlight the effectiveness of WiCLP, which outperforms full fine-tuning of CLIP while being significantly more parameter-efficient. While fine-tuning the last layers of CLIP combined with WiCLP achieves slightly better performance than using WiCLP alone, it requires optimizing a much larger number of parameters. Given this trade-off, we prioritize WiCLP alone to minimize the number of parameters while achieving substantial compositional performance improvements. Additionally, keeping the original CLIP unchanged makes our approach more suitable for Switch-Off functionality, allowing the module to be easily enabled or disabled as needed.
Additionally, we conducted an ablation study on various projection layer architectures, ranging from simple designs to more advanced, parameter-heavy transformer-based models. For detailed descriptions of these architectures, refer to Appendix C. Our evaluation across all attribute binding categories of T2I-CompBench, as shown in Figure 10, reveals that CLP and WiCLP consistently achieve the highest performance gains while remaining both parameter- and speed-efficient. For further results and analysis, please refer to Appendix C.
@@ 6 Conclusion
We analyze error sources in text-to-image models for generating images from compositional prompts, identifying (i) erroneous attention contributions in CLIP token embeddings and (ii) the CLIP text-encoder’s sub-optimal alignment with the UNet. Based on these insights, we propose WiCLP, a simple yet strong baseline that fine-tunes a linear projection on CLIP’s representation space. WiCLP though inherently simple and parameter efficient, outperforms existing methods on compositional image generation benchmarks and maintains a low FID score on a broader range of clean prompts.
Limitations
In this paper, we have conducted a comprehensive analysis of one of the primary reasons why Stable Diffusion struggles with generating compositional attribute binding prompts and proposed a lightweight, efficient method to address this challenge. While our approach demonstrates promising results, there remains substantial room for improvement in this area. Our method primarily targets the attribute binding aspect of compositionality, leaving other critical categories, such as spatial relationships (e.g., "a book to the left of a pen"), numeracy (e.g., "four books"), and others, less explored. Investigating the underlying causes of these issues is crucial for advancing the field further.
Moreover, the reliance on CLIP—particularly the CLIP score—as a metric for recognizing and evaluating compositionality poses its own limitations. CLIP, in its current form, does not perform optimally for such tasks. A promising direction for future research would be to first improve CLIP’s ability to handle compositionality effectively and then adapt this enhanced version of CLIP for Stable Diffusion. This could pave the way for more robust and accurate text-to-image generation models.